# An approach based on metaheuristic algorithms to the timetabling problem in deregulated railway markets

### v1

@@ Abstract
The train timetabling problem in liberalized railway markets represents a challenge to the coordination between infrastructure managers and railway undertakings. Efficient scheduling is critical in maximizing infrastructure capacity and utilization while adhering as closely as possible to the requests of railway undertakings. These objectives ultimately contribute to maximizing the infrastructure manager’s revenues. This paper sets out a modular simulation framework to reproduce the dynamics of deregulated railway systems. Ten metaheuristic algorithms using the MEALPY Python library are then evaluated in order to optimize train schedules in the liberalized Spanish railway market. The results show that the Genetic Algorithm outperforms others in revenue optimization, convergence speed, and schedule adherence. Alternatives, such as Particle Swarm Optimization and Ant Colony Optimization Continuous, show slower convergence and higher variability. The results emphasize the trade-off between scheduling more trains and adhering to requested times, providing insights into solving complex scheduling problems in deregulated railway systems.
@@
[…]
journal: Integrated Computer-Aided Engineering
, and
@@ 1 Introduction
In recent decades, the railway system in European Union countries has undergone major evolution [1], and the main challenge of all the changes the sector is facing is the establishment of a Single European Railway Area (SERA) [2]. Achieving this goal will have two important consequences: on the one hand, it will open up the passenger rail market to new operating Railway Undertakings (RUs). This will lead to an increase in the supply available to users as well as to more competitive prices [3]. On the other hand, a second consequence of the creation of SERA will allow railway operators to provide passenger services in every Member State [3]. The gradual transition towards a liberalized European rail market affects the market structure in the Member States [1]. Currently, there is a vertical separation in railway markets [4]. This structure defines three clearly differentiated players or roles in the market: the Infrastructure Manager (IM), the RUs and the demand [5]. The IM is the entity that, through a contract with the owner of the railway infrastructure, obtains the rights to manage it. Some of its functions include the capacity allocated to the different RUs, track access charging, or the infrastructure maintenance, among others [6]. The RUs are the companies offering rail services. In a liberalized railway market, there is more than one RU and they compete with each other in two ways: firstly, they compete for the infrastructure managed by the IM and, secondly, they compete on price to attract passengers, the third main actor in the market, corresponding to the demand [7]. In this current context, it is crucial to develop advanced tools to solve the complex problems that arise in this dynamic and competitive environment. These include the allocation of time slots to RUs [5], the establishment of timetables, the setting of prices [8], among others. This proposal specifically focuses on solving the timetabling problem in a liberalized railway market.
@@ 1.1 Contributions
This paper focuses on the railway timetabling problem in liberalized railway markets. Specifically, the Spanish railway market has been used as a test case for the computational experiments presented in this work. The main contributions of this paper are the following:
  • Problem modeling: The article introduces a dynamic model tailored to the timetabling problem in a liberalized railway market, addressing the interplay between competition and cooperation. The methodology rigorously incorporates the relationships between infrastructure managers (IM) and railway undertakings (RUs). This analytical tool enables the exploration of various scenarios and policies, forming the groundwork for automated, data-driven decision making in resource allocation within competitive and dynamic market conditions.
  • Solution approach: The study applies metaheuristic and bio-inspired algorithms to tackle the time-tabling problem in a liberalized railway market. The authors systematically adapt these algorithms, known for their success in other scheduling domains, ensuring replicability and consistency. By utilizing MEALPY library [9, 10], the study provides a comprehensive computational framework for addressing complex optimization challenges.
  • Analysis and comparison: The study evaluates the applicability of metaheuristic algorithms to the timetabling problem in liberalized railway markets, analyzing solution quality, run time, convergence, and hyperparameter sensitivity. The findings validate their suitability for this context and highlight the comparative strengths of different algorithmic approaches, enriching the methodological toolkit for scheduling and optimization research.
@@ 1.2 Structure of the paper
The paper is structured as follows: Section 2 reviews metaheuristic algorithms in scheduling, and recent contributions in transport and railway planning. Section 3 formulates the mathematical model for the timetabling problem. Section 4 analyzes the results of the applied algorithms. Finally, Section 5 presents conclusions and future lines of research.
@@ 2 Background
A metaheuristic algorithm is a high-level algorithmic framework that provides a set of strategies for developing heuristic optimization algorithms independently of the problem to be solved [11]. To develop this type of algorithm, the scientific community has been inspired by a multitude of sources of knowledge such as the theory of evolution, biological systems, the behavior of swarms of species, etc [12]. Thanks to this inspiration, a large number of metaheuristic algorithms have appeared which, depending on their source of inspiration, have been classified as evolutionary algorithms, physics-based algorithms, bio-inspired, etc [13, 14]. These algorithms can obtain good solutions to complex problems in a shorter time than exact algorithms, and are thus very useful for solving optimization problems that require good solutions to be obtained in a short time.
2.1 Metaheuristics in scheduling and planning problems
@@
This section explores the application of metaheuristic algorithms to the solution of scheduling and planning problems [15]. Scheduling and planning problems appear frequently in many application domains, such as process planning and scheduling [16]. More specifically, two common problems in the field of production systems are job shop scheduling and flow shop scheduling [17]. One recent review paper is [18], in which the authors analyze more than one hundred papers where metaheuristic algorithms are used to solve hybrid flow shop scheduling problems. In electrical engineering and power systems, metaheuristic algorithms have also been applied to many problems. One of these is the well-known Transmission Network Expansion Planning Problem [19]. In this problem, the objective is to determine the necessary expansions in the electricity transmission network to meet future energy demand in a reliable, economic and safe manner [20]. To solve this problem, metaheuristic algorithms such as Tabu Search (TS) [21] or Particle Swarm Optimization (PSO) [22] have been successfully applied. The good performance of metaheuristic algorithms on this problem is such that recent work has compared the performance of these methods with that of mathematical methods [23]. Transport is another area of application where many planning and scheduling problems arise. For this reason, it is common to find in the literature many studies in which metaheuristic algorithms are applied to solve this type of problem, such as traffic prediction [24]. Similarly, optimization of signalling is another hot optimization problem in the literature as a method for responding to demand in urban areas. In [25], the authors addressed the two problems mentioned above using metaheuristic algorithms such as Hill Climbing (HC), Simulated Annealing (SA) or TS among others, analyzing the performance of all of them. With respect to liberalized railway markets, the work of [26] is a first approach to modeling intelligent negotiation between the IM and the RUs in a liberalized railway market. To do this, a multi-agent reinforcement learning framework based on a Q-Learning algorithm is proposed. The simulation results provided by these authors shows their approach enhances the intelligent behavior of the actors involved in pursuing their objectives. Then, in [27], the PSO algorithm is use to solve the timetabling problem in a liberalized railway market. This paper, unlike the previous one, focuses more on the solution of the timetabling problem than on the bargaining between market players. The simulation experiments carried out by the authors allow them to determine that PSO is capable of generating feasible timetables that allow the IM to maximize track access charges as well as capacity use. Game theory is another approach used to solve the railway planning problem. In [28], game theory allows not only the problems of interaction/negotiation between different agents to be solved, but also the behavior of each of these agents, who want to maximize their own objectives or interests, to be identified and understood. Taking this work as a starting point, these same authors developed, in [29], a multi-agent system based on a cooperative game-theoretic non-transferable utility approach, in order to solve the timetabling problem. The results obtained by the authors in a simulated environment of the Istanbul-Ankara high-speed rail network provided a Nash solution with profitable services and an acceptable level of satisfaction for passengers. The problem of timetabling is not a simple one and even less so in a liberalized environment. In [30], a total of seven quality aspects that a timetable should have are identified and a study is carried out in which eight Swedish practitioners were interviewed to find out the current state of practice in that country.
@@ 2.2 Metaheuristic algorithms of study
This section details the metaheuristic algorithms applied to solving the train timetabling problem in a liberalized railway system. For each, their implementation in the Python MEALPY library has been taken as a reference. In summary, Table 1 shows each of the algorithms used, a brief description, and their references, including the applications of each algorithm in transport problems (if known).
@@
Genetic Algorithm (GA) Inspired by natural selection and genetics. Evolves a population through selection, crossover, and mutation. Widely applied in railway scheduling, rolling stock circulation, and timetabling. [31] [32, 33, 34]
Simulated Annealing (SA) Mimics the annealing process in metallurgy. Explores the solution space probabilistically, accepting worse solutions to escape local optima. Used in timetable design, maintenance scheduling, and rolling stock problems. [35, 36] [37, 38, 39]
Ant Colony Optimization Continuous (ACOR) Extends the ACO paradigm to continuous optimization. Simulates the pheromone-laying behavior of ants. Applied in train routing, rolling stock planning, and timetabling. [40] [41, 42, 43]
Particle Swarm Optimization (PSO) Inspired by the social behavior of birds and fish. Optimizes by iteratively updating candidates based on velocity and position. Commonly used in timetable optimization. [44] [27, 45, 46]
Differential Evolution (DE) A population-based algorithm combining differences between individuals to evolve solutions. Recently applied to train delay scheduling and timetable optimization. [47, 48] [49, 50]
Artificial Bee Colony (ABC) Simulates the foraging behavior of honey bees. Balances exploration and exploitation through employed, onlooker, and scout bees. Recently applied in timetable design. [51, 52] [53, 54]
Covariance Matrix Adaptation Evolution Strategy (CMA-ES) Adapts the covariance matrix to guide sampling. Limited transport applications, but used in railway real-time scheduling. [55] [56]
Grey Wolf Optimizer (GWO) Mimics the hierarchy and hunting behavior of grey wolves. Uses alpha, beta, and delta wolves to lead the search process. Applied in railway scheduling. [57] [58, 59]
Whale Optimization Algorithm (WOA) Inspired by the bubble-net feeding behavior of whales. Alternates between encircling prey and random search. Limited transport applications, included due to its recent popularity. [60] -
Hybrid Grey Wolf - Whale Optimization Algorithm (GWO-WOA) Combines strengths of GWO and WOA. Leverages GWO leadership and WOA encircling to enhance performance. Selected to compare hybrid and standalone algorithm efficiencies. [61] -
Figure 1: Relationship between the RUs and IM

### v2

# An approach to the timetabling problem in deregulated railway markets based on metaheuristic algorithms

@@ Abstract
The train timetabling problem in liberalized railway markets represents a challenge to the coordination between infrastructure managers and railway undertakings. Efficient scheduling is critical to maximizing infrastructure capacity and utilization while adhering as closely as possible to the requests of railway undertakings. These objectives ultimately contribute to maximizing the infrastructure manager’s revenues. This paper sets out a modular simulation framework to reproduce the dynamics of deregulated railway systems. Ten metaheuristic algorithms using the MEALPY Python library are then evaluated in order to optimize train schedules in the liberalized Spanish railway market. In addition, an analysis of the scalability of the problem has been carried out by comparing the results with those obtained with a classical mathematical model such as SCIP in Pyomo. The results show that the Genetic Algorithm outperforms others in revenue optimization, convergence speed, and schedule adherence. Alternatives, such as Particle Swarm Optimization and Ant Colony Optimization Continuous, show slower convergence and higher variability. The results emphasize the trade-off between scheduling more trains and adhering to requested times, providing insights into solving complex scheduling problems in deregulated railway systems.
@@
Railway Transportation, Optimization, Metaheuristics,Timetabling, Scheduling
@@ 1 Introduction
In recent decades, the railway system in European Union countries has undergone major evolution Ait2021 , and the main challenge of all the changes the sector is facing is the establishment of a Single European Railway Area (SERA) COM2011 . Achieving this goal will have two important consequences: on the one hand, it will open up the passenger rail market to new operating Railway Undertakings (RUs). This will lead both to an increase in the supply available to users and to more competitive prices Caramello2017 . On the other hand, a second consequence of the creation of SERA will allow railway operators to provide passenger services in every member state Caramello2017 . Currently, projects such as the Timetabling and Capacity Redesign for Smart Capacity Management (TTR), seek to improve the flexibility, efficiency and cost-effectiveness of the market to meet today’s needs  [^1] Forum Train Europe FTE, Introduction to TTR: https://www.forumtraineurope.eu/services/ttr . For example, in Switzerland, capacity planning under this program is based on the concepts of network usage (NNK) and network usage plan (NNP), aligning the capacity assignment strategies with the existing national tools available and the neighboring countries. The gradual transition towards a liberalized European rail market affects the market structure in the member states Ait2021 . Currently, there is a vertical separation in railway markets Bouf05 . This structure defines three clearly differentiated players or roles in the market: the Infrastructure Manager (IM), the RUs and the demand Besinovic2024 . The IM is the entity that, through a contract with the owner of the railway infrastructure, obtains the rights to manage it. Some of its functions include the capacity allocated to the different RUs, track access charging, or infrastructure maintenance, among others Pena2015 . The RUs are the companies offering rail services. In a liberalized railway market, there is more than one RU and they compete with each other in two ways: firstly, they compete for the infrastructure managed by the IM and, secondly, they compete on price to attract passengers, the third main actor in the market, corresponding to the demand Link04 . This contrasts with the classic monopolistic framework, where a single RU both plans and operates services under unified control, greatly simplifying the timetabling process. Whereas traditional timetabling assumes centralized decision-making, liberalized markets require coordination, strategic bidding for slots and dynamic pricing considerations. In this current context, it is crucial to develop advanced tools to solve the complex problems that arise in this dynamic and competitive environment. These include the allocation of time slots to RUs Besinovic2024 , the establishment of timetables, the setting of prices Smoliner2018 , among others. This proposal specifically focuses on solving the timetabling problem in a liberalized railway market.
@@ 1.1 Contributions
This paper focuses on the railway timetabling problem in liberalized railway markets. Specifically, the Spanish railway market has been used as a test case for the computational experiments presented in this study. The main contributions of this paper are the following:
  • A novel RU competitive timetabling model. This study introduces a dynamic mathematical formulation that models the interactions between IM and RUs in a realistic and adaptable way. Its main innovation lies in the development of an analytical tool capable of simulating multiple scenarios and policy options, paving the way for automated, data-driven decision making in the efficient allocation of resources within highly competitive and dynamic railway markets.
  • A heuristic for scheduling conflicting services: The paper introduces a deterministic greedy algorithm that converts the continuous departure‐time vector produced by the metaheuristics into a discrete, conflict‐free service schedule. The heuristic first identifies and schedules all services without conflicts, then iteratively selects the highest‐revenue service among those in conflict, adds it to the schedule, and removes its conflicting services from further consideration. By doing so, it reduces the combinatorial complexity, delivering consistent, high‐quality schedules even for large service sets.
  • Customized bio-inspired optimization framework. This study advances the state of the art by leveraging the open-source MEALPY library to apply metaheuristic and bio-inspired algorithms to the complex timetabling problem in a liberalized railway market—an area where such techniques have been under-explored. Unlike prior work, which often lacks transparency or adaptability, this research systematically adapts and integrates these algorithms—proven effective in other scheduling domains—into a replicable and consistent frameworkwhich is publicly available in the following repository [^2] https://github.com/DavidMunozValero/GSA_M to ensure that the results are reproducible. By building on MEALPY Vanthieu2023 ; Vanthieu2023b , the study not only demonstrates the applicability of these methods to railway operations, but also introduces a robust, open computational environment that facilitates experimentation and benchmarking in complex optimization scenarios.
  • Metaheuristic comparison under uniform conditions. This study makes a novel contribution by systematically benchmarking a diverse set of metaheuristic algorithms—from classical to more recent approaches—under identical experimental conditions. Unlike previous research, which often evaluates algorithms in isolation or under varying setups, this work provides a unified and rigorous comparison framework. It analyzes not only solution quality and runtime, but also convergence behavior and sensitivity to hyperparameters, offering a comprehensive view of algorithmic performance. Furthermore, a comparison with a classical mathematical approach has been carried out by using Pyomo hart2011pyomo and the Solving Constraint Integer Programs (SCIP) solver scip . The findings reveal meaningful trade-offs between quality, efficiency, and robustness, offering practical guidance for algorithm selection in railway timetabling. The direct, side-by-side comparison of these algorithms in this specific context represents a significant advance over the existing literature.
@@ 1.2 Structure of the paper
The paper is structured as follows: Section 2 reviews metaheuristic algorithms in scheduling, and recent contributions in transport and railway planning. Section 3 formulates the mathematical model for the timetabling problem. Section 4 analyzes the results of the algorithms applied. Finally, Section 5 presents conclusions and future lines of research.
@@ 2 Background
A metaheuristic algorithm is a high-level algorithmic framework that provides a set of strategies for developing heuristic optimization algorithms independently of the problem to be solved Sorensen13 . To develop this type of algorithm, the scientific community has been inspired by a multitude of sources of knowledge, such as the theory of evolution, biological systems, the behavior of swarms of species, etc delSer19 . Thanks to this, a large number of metaheuristic algorithms have appeared which, depending on their source of inspiration, have been classified as evolutionary algorithms, physics-based algorithms, bio-inspired, etc Talbi02 ; Rajwar23 . These algorithms can return good solutions to complex problems in a shorter time than exact algorithms, and are thus very useful for solving optimization problems that require good solutions in a short time.
The following sections contain a literature review regarding the railway timetabling approaches and the applied metaheuristics algorithms.
2.1 Railway timetabling approaches
Metaheuristic algorithms have been successfully applied to solving scheduling and planning problems in many fields in the literature Zarandi20 . These applications range from production systems Neufeld23 to process planning Li15 and electrical power systems Morquecho24 . This remarkable success has led the transport community to focus on these methods as an alternative to the exact mathematical methods commonly used in this field of application. Thus, this section explores the application and approaches used in the literature to the solution of the railway timetabling problem.
Railway timetabling is a hot topic in the transport and operations research communities. Many contributions and breakthroughs have been made in this field in the last fifteen years. Traditionally, optimal railway timetables were calculated to minimize passenger waiting times. However, this waiting-time-focused approach neglected to deal with train delays, resulting in timetables that were not as efficient as they should have been. In Liebchen2010 , an integer-linear programming model was proposed to compute delay resistant railway timetables to fill this research gap. Derived from train delays, canceled connections are another factor affecting the efficiency of timetables. Based on the previous work, Corman2012 developed a multi-objective optimization approach to compute optimal timetables minimizing train delays and missed connections. This approach allows it to be considered which connection to keep or drop in order to reach a compromise solution. During these years, many optimization models emerged for timetable optimization and real-time re-planning of railway services. A wide review of all these approaches can be found in Cacchiani2014 .
Later, dynamic demand considerations were introduced in the proposed formulations to solve the railway timetabling problem. For example, Barrena2014 proposed two mathematical programming formulations considering a dynamic demand pattern in a single-line rail rapid transit system. Furthermore, the authors proposed a heuristic tailored algorithm to solve large instances of this problem, improving the results obtained by other solvers like CPLEX. The same authors, in Barrena2014_2 , developed three mathematical programming formulations for the train timetabling problem with dynamic demand, and the problem was solved by means of a branch-and-cut approach.
Considering demand in a dynamic way to solve the timetabling problem led to the formulation of the passenger-centric timetabling problem Robenek2016 . In contrast to traditional approaches where the aim was to minimize waiting times and delays, the passenger-centric timetabling problem aims to maximize customer satisfaction in timetable design. However, the methods for its resolution were still mixed integer linear programming models. Then, more optimization models based on this approach were proposed. A good example is the work of Binder2017 , where the authors proposed a multi-objective optimization model in order to maximize passenger satisfaction, and minimize the operational costs and the deviation from an undisrupted timetable. Other recent work is Cacchiani2020 , where a mixed-integer linear programming model was introduced to solve the timetabling and stop planning problem considering an uncertain demand.
@@ 2.1 Railway timetabling approaches
With respect to liberalized railway markets, the work of Wong10 was a first approach to modeling intelligent negotiation between the IM and the RUs in a liberalized railway market. A multi-agent reinforcement learning framework based on a Q-Learning algorithm was proposed for this purpose. The simulation results provided by these authors showed that their approach enhances the intelligent behavior of the actors involved in pursuing their objectives. Then, in Ho12 , the Particle Swarm Optimization (PSO) algorithm was used to solve the timetabling problem in a liberalized railway market. This paper, unlike the previous one, focused more on the solution of the timetabling problem than on the bargaining between market players. The simulation experiments carried out by the authors allowed them to determine that PSO was capable of generating feasible timetables that allowed the IM to maximize both track access charges and capacity use. Game theory is another approach used to solve the railway planning problem. In Shakibaei19 , game theory allowed not only the problems of interaction/negotiation between different agents to be solved, but also the behavior of each of these agents, who want to maximize their own objectives or interests, to be identified and understood. Taking this paper as a starting point, these same authors developed, in Shakibaei21 , a multi-agent system based on a cooperative game-theoretic non-transferable utility approach, in order to solve the timetabling problem. The results obtained by the authors in a simulated environment of the Istanbul-Ankara high-speed rail network provided a Nash solution with profitable services and an acceptable level of passenger satisfaction. The problem of timetabling is not a simple one, and even less so in a liberalized environment. In Gestrelius20 , a total of seven quality aspects that a timetable should have are identified and a study is carried out in which eight Swedish practitioners were interviewed to find out the current state of practice in that country.
Finally, there are different recent studies for train scheduling in complex railway stations, tackling aspects such as track assignment, coordination, and conflict minimization. For example, Zhong24 proposed a parallel optimization method for the simultaneous scheduling of trains and shunting operations at complex high-speed railway stations. In addition, Liu23 developed a control strategy for the stable formation of high-speed virtually coupled trains under disturbances and delays, ensuring synchronized movements and reduced headway conflicts— an approach that can further enhance overall scheduling robustness and line capacity use.
In summary, the train timetabling problem has traditionally been tackled using mathematical programming methods like integer and mixed-integer linear programming. While some studies have applied metaheuristic algorithms tailored to the problem, there is still a gap in understanding their usefulness for this problem. Metaheuristics offer advantages such as being derivative-free, easy to implement, and capable of providing quick, approximate solutions— especially valuable in complex or time-sensitive scenarios. This paper attempts to address this research gap by designing a framework for using any metaheuristic algorithm (not necessarily designed purposely for this problem). Furthermore, this work aims to address that gap by systematically evaluating ten metaheuristic algorithms to establish a baseline for their potential and suitability in solving the train timetabling problem. Then, the performance of metaheuristic algorithms is compared with exact mathematical methods in order to provide a benchmark for the timetabling problem and understand the main advantages of the proposed approach.
@@ 2.2 Metaheuristic algorithms of study
This section details the metaheuristic algorithms applied to solving the train timetabling problem in a liberalized railway system. For each, their implementation in the Python MEALPY library has been taken as a reference, since it is known for its comprehensiveness and citation impact. A representative set of ten metaheuristic algorithms was selected based on three criteria: (1) Historical coverage, including both classical and recent algorithms, from Simulated Annealing in 1985 to a Hybrid Grey Wolf–Whale Optimization Algorithm in 2022; (2) Diversity of types, covering single-individual algorithms (such as SA), those population-based, including evolutionary (GA, DE, CMA-ES), bio-inspired (ACOR, PSO, ABC, GWO, WOA) and hybrid algorithms, like the GWO-WOA; (3) Performance, with all algorithms recognized for their effectiveness and literature relevance, and many of them (GA, PSO, DE and CMA-ES) provide the basis of algorithms that have frequently won major international optimization competitions, such as IEEE CEC Qiao24 and GECCO [^3] https://www.gecad.isep.ipp.pt/ERM-competitions/2024-2/.
In summary, Table 1 shows each of the algorithms used, with a brief description, and their references, including the applications of each algorithm in transport problems (if known).
@@
Genetic Algorithm (GA) Inspired by natural selection and genetics. Evolves a population through selection, crossover, and mutation. Widely applied in railway scheduling, rolling stock circulation, and timetabling. Holland75 Canca18 ; Pan20 ; Han21
Simulated Annealing (SA) Mimics the annealing process in metallurgy. Explores the solution space probabilistically, accepting worse solutions to escape local optima. Used in timetable design, maintenance scheduling, and rolling stock problems. Cerny85 ; Kirkpatrick85 Yue17 ; Robenek18 ; Lin19
Ant Colony Optimization Continuous (ACOR) Extends the ACO paradigm to continuous optimization. Simulates the pheromone-laying behavior of ants. Applied in train routing, rolling stock planning, and timetabling. Dorigo92 Tsuji12 ; Sama16 ; Coviello23
Particle Swarm Optimization (PSO) Inspired by the social behavior of birds and fish. Optimizes by iteratively updating candidates based on velocity and position. Commonly used in timetable optimization. Eberhart95 Ho12 ; Jamili12 ; Wu16
Differential Evolution (DE) A population-based algorithm combining differences between individuals to evolve solutions. Recently applied to train delay scheduling and timetable optimization. Storn95 ; Storn97 Zhong13 ; Song23
Artificial Bee Colony (ABC) Simulates the foraging behavior of honey bees. Balances exploration and exploitation through employed, onlooker, and scout bees. Recently applied in timetable design. Karaboga05 ; Karaboga07 Chen22 ; Wang24
Covariance Matrix Adaptation Evolution Strategy (CMA-ES) Adapts the covariance matrix to guide sampling. Limited transport applications, but used in railway real-time scheduling. Auger11 Prasad20
Grey Wolf Optimizer (GWO) Mimics the hierarchy and hunting behavior of gray wolves. Uses alpha, beta, and delta wolves to lead the search process. Applied in railway scheduling. Mirjalili14 Yu23 ; Zhang24
Whale Optimization Algorithm (WOA) Inspired by the bubble-net feeding behavior of whales. Alternates between encircling prey and random search. Limited transport applications, included due to its recent popularity. Mirjalili16 -
Hybrid Grey Wolf - Whale Optimization Algorithm (GWO-WOA) Combines strengths of GWO and WOA. Leverages GWO leadership and WOA encircling to enhance performance. Selected to compare hybrid and standalone algorithm efficiencies. Obadina22 -